OpenStudy (anonymous):

the base of a triangle is 7 cm more than its height. if the area is 294 square centimeter, find the base and the height of the triangle.

OpenStudy (anonymous):

Thank you. :)

OpenStudy (anonymous):

your welcome..

OpenStudy (anonymous):



OpenStudy (anonymous):

what is the representation for that?

OpenStudy (anonymous):

\[294={1\over2}(7+h)(h)\]

OpenStudy (anonymous):

Can you solve for h?

OpenStudy (anonymous):

294=1/2 (7h+h squared)

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Now, bring 1/2 over to the other side.

OpenStudy (anonymous):

so it become 294-1/2= 7x+x squared

OpenStudy (anonymous):

No. when you bring over a fraction, it becomes inverted/multiplied

OpenStudy (anonymous):

\[294 \times{2 \over 1} =7h+h^2\]

OpenStudy (anonymous):

what happened to the fraction?

OpenStudy (anonymous):

hmm... Let me explain it in another way.

OpenStudy (anonymous):

so it become 588? right?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

So now, \[588=7h+h^2\]

OpenStudy (anonymous):

because... \[{588 \over 1}=588\]

OpenStudy (anonymous):

You still following?

OpenStudy (anonymous):

then it become h squared+7h - 588? right?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Yes. :)

OpenStudy (anonymous):

now you can use the quadratic equation to solve.

OpenStudy (anonymous):

you know how to, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Good. :)

OpenStudy (anonymous):

how can i saw our conversation before?

OpenStudy (anonymous):

is the answer 24 and 28?

OpenStudy (anonymous):

\[{-b \pm \sqrt{b^2-4ac}} \over 2a\]

OpenStudy (anonymous):

do only the plus.

OpenStudy (anonymous):

it's not 24 but -28 is correct. This one is the one you don't need to use.

OpenStudy (anonymous):

its 21 and 28

OpenStudy (anonymous):

yes. The answer is 21 since you only want the positive integer.

OpenStudy (anonymous):

so h=21. What is the base? 7+h=?

OpenStudy (anonymous):

21+7 = 28

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so the base is 28, the height is 21.